On January 1, 2021, the company obtained a $3 million loan with a 12% interest rate. The building was completed on September 30, 2022. Expenditures on the project were as follows:

January 1, 2021 $ 1,170,000
March 1, 2021 990,000
June 30, 2021 110,000
October 1, 2021 730,000
January 31, 2022 855,000
April 30, 2022 1,170,000
August 31, 2022 2,070,000


On January 1, 2021, the company obtained a $3 million construction loan with a 12% interest rate. Assume the $3 million loan is not specifically tied to construction of the building. The loan was outstanding all of 2021 and 2022. The company’s other interest-bearing debt included two long-term notes of $5,300,000 and $7,300,000 with interest rates of 5% and 7%, respectively. Both notes were outstanding during all of 2021 and 2022. Interest is paid annually on all debt. The company’s fiscal year-end is December 31.

Required:
1. Calculate the amount of interest that Mason should capitalize in 2021 and 2022 using the weighted-average method. Calculate the interest expense for 2021 and 2022.
2. What is the total cost of the building?
3. Calculate the amount of interest expense that will appear in the 2021 and 2022 income statements.
Answer #1
Interest to be capitalized in 2020:
Weighted average accumulated expenditure:
Amount Months Weighted
Average
expenses
Jan 1.Expenses 1170000 12 (1170000*12/12)=1170000
Mar 1.Expenses 990000 10 (990000*10/12)=825000
June 30.Expenses 110000 6 (110000*6/12)=55000
Oct 1.Expenses 730000 3 (730000*3/12)=182500
Total 2232500
Month is calculated from the date of expenditure incurred till Dec 31.
Weighted average interest rate
Amount
of loan
Interest % Interest
12% loan 3000000 12% 360000
5% note 5300000 5% 265000
7% note 7300000 7% 511000
15600000 1136000
Weighted average interest rate=Total interest expenses/Total amount of loan=1136000/15600000=0.0728=7.28%
Avoidable interest=Weighted average accumulated expenditure*Weighted average interest rate=2232500*7.28%=$ 162526
Actual interest paid=$ 1136000
Since total avoidable interest is less than actual interest,Interest to be capitalized=Avoidable interest=$ 162526
Interest to be capitalized in 2021:
Weighted average accumulated expenditure:
Amount Months Weighted
Average
expenses
Jan 31.Expenses 855000 8 (855000*8/9)=760000
Apr 30.Expenses 1170000 5 (1170000*5/9)=650000
Aug 31.Expenses 2070000 1 (2070000*1/6)=345000
Total 1755000
Construction completed on Sep 30.Hence, total months=9 months
Month is calculated from the date of expenditure incurred till Sep 30.
Weighted average interest rate:
Amount
of loan
Interest % Interest till sep 30.
12% loan 3000000 12% 270000
5% note 5300000 5% 198750
7% note 7300000 7% 383250
15600000 852000
Weighted average interest rate=Total interest expenses/Total amount of loan=852000/15600000=0.0546=5.46%
Avoidable interest=Weighted average accumulated expenditure*Weighted average interest rate=1755000*5.46%=$ 95823
Actual interest paid=$ 852000
Since total avoidable interest is less than actual interest,Interest to be capitalized=Avoidable interest=$ 95823
2 Total cost of building:
$
Total expenditures on the project
(1170000+990000+110000+730000+855000+1170000+2070000) 7095000
Interest capitalized
2020 162526
2021 95823
Total cost of building 7353349
3 Interest expense to be recognized in income statement=Actual interest-Interest capitalized
2020 Interest expense to be recognized in income statement=1136000-162526=$ 973474
2021 Interest expense to be recognized in income statement=852000-95823=$ 756177